OpenStudy (anonymous):

Find two real numbers whose sum is 24 and whose product is minimum.

OpenStudy (mathstudent55):

Do you mean a mimimum or a maximum?

OpenStudy (anonymous):

minimum :)

OpenStudy (mathstudent55):

I don't think there can be a minimum in this probelm.

OpenStudy (anonymous):

Hmm, i see. That's what i've been thinking of. Well, thank yoooou so much for the help @mathstudent55 ^_^v

OpenStudy (mathstudent55):

This problem will give you a function whose graph is a parabola that opens downward, so it has a maximum value, but it extends downard for ever. There is no minimum value.

OpenStudy (anonymous):

Hmm. Okaaaay :) thank you for your explaination @mathstudent55 and thank you for your help so much! ^_^ you really helped alot. Arigato gozaimas. ^_^v

OpenStudy (mathstudent55):

Let on number by x The other number is 24 - x Their product is x(24 - x) = 24x - x^2 y = -x^2 + 24x has a vertex at x = 24-2 = 12 The _maximum_ product is when both numbers are 12.